OpenStudy (anonymous):

Solve for x: 6x + 12 = 4(x + 6) A) x = 6 B) x = -6 C) x = 9 D) x = -9

OpenStudy (anonymous):

2 x-12 = 0 6 (x+2) = 4 (x+6) 6 x+12 = 4 x+24

OpenStudy (anonymous):

so therefore the ans is x = 6

OpenStudy (anonymous):

u can check by solving it

OpenStudy (australopithecus):

6x + 12 = 4(x + 6) 6x + 12 = 4x + 6(4) 6x - 4x + 12 - 24 = 0 2x - 12 = 0 2x = 12 x = 12/2 x = 6

OpenStudy (anonymous):

oh that makes alot of sense now! thanks so much!!